# RAG vs. Long Context: Examining Frontier Large Language Models for Environmental Review Document Comprehension

### v1

@@ Abstract
Large Language Models (LLMs) have been applied to many research problems across various domains. One of the applications of LLMs is providing question-answering systems that cater to users from different fields. The effectiveness of LLM-based question-answering systems has already been established at an acceptable level for users posing questions in popular and public domains such as trivia and literature. However, it has not often been established in niche domains that traditionally require specialized expertise. To this end, we construct the NEPAQuAD1.0 benchmark to evaluate the performance of three frontier LLMs – Claude Sonnet, Gemini, and GPT-4 – when answering questions originating from Environmental Impact Statements prepared by U.S. federal government agencies in accordance with the National Environmental Environmental Act (NEPA). We specifically measure the ability of LLMs to understand the nuances of legal, technical, and compliance-related information present in NEPA documents in different contextual scenarios. For example, we test the LLMs’ internal prior NEPA knowledge by providing questions without any context, as well as assess how LLMs synthesize the contextual information present in long NEPA documents to facilitate the question/answering task. We compare the performance of the long context LLMs and RAG powered models in handling different types of questions (e.g., problem-solving, divergent). Our results suggest that RAG powered models significantly outperform the long context models in the answer accuracy regardless of the choice of the frontier LLM. Our further analysis reveals that many models perform better answering closed questions than divergent and problem-solving questions.
@@ 1 Introduction
As Large Language Models (LLMs) become increasingly commonplace, researchers have discovered that these models are useful for many tasks beyond text generation. Specifically, LLMs have shown potential utility in niche domains (like science) that would traditionally require specialized expertise, both in a pure text setting Horawalavithana et al. (2022); Munikoti et al. (2024) and by incorporating data of various modalities Dollar et al. (2022); Horawalavithana et al. (2023). Recent work has been done to evaluate these models Acharya et al. (2023); Munikoti et al. (2023); Cai et al. (2024) and to assess their uncertainty Wagle et al. (2024). Despite extensive research, constructing LLMs for answering domain-specific questions has proven challenging Kasneci et al. (2023).
@@
In this work, we focus on assessing the long-context LLMs in the environmental reviews conducted under the National Environmental Policy Act (NEPA) [^1] https://www.epa.gov/nepa. NEPA is a U.S. environmental law designed to protect the environment. An environmental impact statement (EIS) is required by Section 102(2)(C) of NEPA for any proposed major federal action significantly affecting the quality of the environment. An EIS is a detailed document that describes a proposed action, alternatives to the proposed action, and potential effects of the proposed action and alternatives on the environment. An EIS contains information about environmental permitting and policy decisions and considers a range of reasonable alternatives, analyzes the potential impacts resulting from the proposed action and alternatives, and demonstrates compliance with other applicable environmental laws and executive orders.
@@
In this work, we leverage both long context and Retrieval-Augmented Generation (RAG) Lewis et al. (2021) to develop LLM capability for question-answering over EIS documents (Figure 1). We select frontier LLMs for our experiments: Claude Sonnet Team and Collaborators , Gemini Team and Collaborators (2024), and GPT-4 OpenAI (2024). To assess the efficacy of our proposed RAG model compared to other context-augmentation strategies, we also conduct rigorous experiments evaluating LLMs with various types of contexts for NEPA documents. To evaluate our approach, we establish a benchmark using ground truth triples of questions, answers, and corresponding contexts, generated through a semi-supervised method employing GPT-4. Overall, we make the following contributions:
@@ 2.1 Gold Passage Selection
Excerpt Selection For each of nine selected documents, we attempt to select excerpts that have important content of each document. Again, the default approach of randomly extracting excerpts poses the risk of resulting in low-quality excerpts, such as parts of appendix or images’ captions. To avoid this risk, NEPA experts manually select excerpts from the documents. They divided each document into three sections: beginning, middle, and end, and then selected two, six, and six excerpts from each of these sections respectively, for a total of 10 excerpts from each document. We use these excerpts as the ground-truth context, called gold passages, for question benchmark generation.
@@ 2.2 Question Type Selection
Once we identified the gold passages, NEPA experts select the type of questions to include in the benchmark. We started with a list of 15 types of questions [^3] https://tinyurl.com/3akej8ct, and eventually narrowed it down to 10 types of questions after extensive discussions. These types are shown in Table 1. In addition to selecting the question types, the NEPA experts also created sample questions for each category for the EIS document domain. For a more detailed description of the question types, as well as example questions, please see Appendices A and B.
@@ 2.4 Ground-truth Benchmark Generation
Automated Generation We selected GPT-4 as the generative model for this task, as GPT-4 has been used for generating questions and answers for documents in other domains such as agriculture Balaguer et al. (2024). Specifically, we use GPT-4 Azure OpenAI service with default setting to execute the question generation prompts. For each document in our nine selected documents, we have 10 gold passages, results in 90 gold passages in total. For each gold passage, we generated 10 sets of questions for each of the 10 question types. We then filtered the generations for incorrect formats. Overall, we generated a benchmark of 1599 ground-truth triples of question-answer-proof over the nine selected EIS documents. We made the benchmark publicly available  (PolicyAI, 2024).
Quality Check In order to judge the generated benchmark’s correctness, we randomly select 100 sample questions for validation. The validation was done by the NEPA experts (two co-authors). Each of them independently went through the sample questions, checking both the correctness of question type (i.e. whether the generated question was the same type of question as requested) and the correctness of the answer. For each question, if either of the evaluators marked the question or answer as correct, we labelled them as correct. Overall, our generated benchmark achieved $77\%$ of answer correctness and $82\%$ of question type correctness.
@@ 3 Experimental Setup
In this work, we experimented with three frontier LLMs: Claude-3 Sonnet, Gemini, and GPT-4. For the context provided to the model, we had four possible variations: no context, PDF documents, silver passages (RAG setup), or gold passages. The combination of the models and context settings resulted in a total of 12 unique configurations (Figure 1). We explain these configurations in details in Section 3.1 and report the evaluation metrics in Section 3.2.
@@ 3.1 Context Variation
No Context: In this setting, we simply query the models with the questions with no context provided, the same as in other general domains. We do not provide any additional context about the origin of the questions, so the models were expected to answer questions from their existing knowledge. While this strategy can work well with popular domains such as sport or literature, we assume that NEPA domain may be challenging for the LLM models to get the accurate answer. In the other word, this setting can be said to be a test of the LLMs to answer out-of-general-domain questions. As such, this setting is usually expected to return low performance.
Full PDF as Context: In this scenario, in addition to the question, we also provide the model the PDF (text) document from which the context to the question was extracted. Since we do not inform the model which part of the document to look at, the generated responses’ accuracy will heavily depend on the models’ ability to pick the correct context from the very large scale textual information provided. We expect this setting to yield performance better than no context.
RAG Context (Silver Passages): In RAG models, when a question is inputted for LLM generation, the corresponding context is extracted as a relevant passage from a given EIS document. We use the standard RAG setup where we encode both question and retrieved passages with BGE embedding model Xiao et al. (2023). We use the cosine similarity score to assess the similarity between the question and the contexts. The number of top-ranked relevant passages extracted, referred to as top-K silver passages, is set at $k=3$.
Gold Passage as Context: In this configuration, we include the actual context from which the question was generated in the prompt, alongside the question content. While the scenario where users manually identify the correct passage is rare in practice, we simulate this scenario to measure how well LLMs can perform if we were able to extract relevant passages with very high accuracy. We expect this setting to perform the best.
@@ 3.2 Evaluation Metrics
In order to evaluate the performance of the models in different configurations, We compare the answers from the generated responses of the model across these various configurations. Overlap based metrics such as BLEU score Papineni et al. (2002), while used by many prior works, simply measure the syntactic similarity, and as such is not suitable to perform evaluations where semantics is more important. As such, for our work, we use the RAGAs score proposed by Es et al. (2023): the Answer Correctness (called Correctness in this paper).
The Answer Correctness score combined two aspects of factual correctness and semantic correctness for its calculation. While factual correctness captured the correctness at phrase/clause level of input answer, the semantic score is achieved by comparing the similarity between vectors of expected answers and predicted answers by using embedding models. GPT-4 is used in calculating the answer correctness that quantifies the factual overlap between the generated answer and the ground truth answer Es et al. (2023). We use the BGE Xiao et al. (2023) as the embedding model for semantic correctness calculation. We set the weight of factual correctness as 0.75 and the weight of semantic correctness as 0.25 for measuring the Answer Correctness.
@@ 4 Performance Analysis
In this section, we describe the overall performance of the LLMs in the question/answering task evaluated with NEPAQuAD1.0 (as presented in Section 2). First, we compare the performance of three frontier LLMs: Claude-3 Sonnet, Gemini, and GPT-4 across various QA contexts (Section 4.1). Second, we compare the model performance across various question types (Section 4.2). In Section 4.3, we evaluate how the models performing to the questions generated from different parts of the document. Finally, we analyze the performance.
@@ 4.1 Evaluating Different QA Contexts
Table 2 reports the overall performance of the models across various QA contexts used in the evaluation. We observed that for the task with no context, the Gemini model produces the most accurate results by far. However, when PDF documents are provided as context, this trend is reversed, with GPT-4 surpassing Gemini in correctness. Despite that Gemini is able to handle very long contexts (1.5M tokens), it is surprising to see its performance drop when provided with PDF documents as additional contexts. This may be due to the model struggling to reason over the large amount of relevant and irrelevant content in the EIS document.
Overall, RAG models perform better in comparison to the models provided with PDF documents as additional contexts. In RAG setup, The Claude model outperforms both other models in term of correctness, although the scores across the Claude and GPT-4 models are much closer. There is notable increase in Gemini’s performance in the RAG setup when compared with the PDF contexts.
As expected, all models perform best on average when provided with the gold passage in comparison to other context variations. In this scenario, model needs to synthesize information that directly contains the answer to the question posed to the model. Notably, models perform comparably when they are provided with the RAG and gold passage contexts.
Context Claude Gemini GPT-4
None 21.50% 50.16% 20.28%
Complete PDF 23.47% 46.62% 56.40%
Silver Passages 68.74% 57.06% 66.86%
Gold Passage 68.41% 61.81% 67.66%
Table 2: Evaluation on the answer correctness of LLMs over different configurations of context over EIS documents. Silver passages are selected by the RAG model.
@@ 4.2 Evaluating Different Question Types
We analyze the performance of LLMs over different types of questions depicted in Figure 3. When analyzing the results based on the type of questions, we see that all three models have superior performance on closed questions when provided with either silver or gold data as context, while GPT-4 is the only model that performs well on these questions even when provided with just the PDFs as context. For all other categories, both Claude and GPT-4 exhibit similar behavior pattern when provided with none or PDF context, although GPT4’s performance is notably better than Claude’s in almost every category, with this difference particularly noticeable with PDF context.
Interestingly, Gemini performs really well when provided with no context at all, and the performance decreases for all categories except the convergent and recall questions when provided with the PDFs as context. For a majority of the question types, even the silver or gold data is not able to get the performance to the level of no-context, with the closed and recall questions being the exceptions. Overall, RAG models and models with other contexts performed best in answering closed questions and worst in answering divergent and problem-solving questions.
4.3 Evaluating Positional Knowledge
(a) Claude
(b) Gemini
(c) GPT-4
Figure 4: The evaluation results measured by the Answer Correctness scores of each LLM used over different parts with silver passages provided as context
We also analyzed the performance of various question types based on the portion of the document from which they were derived, as shown in Figure 4. Across the models, we observed a general trend where the earlier the source text in the document, the better the performance of the models. A notable exception to this are the problem-solving questions, which perform better when sourced from the latter parts of the document. This pattern holds true for all three models. Additionally, we noticed that divergent questions yield better results when derived from the middle of the document. Overall, all three models exhibit similar or comparable patterns of performance across different document sections and question types. These results suggest that performance of long-context models may vary not only by the position of relevant information, but also due to the type of the question and the amount of reasoning that the model has to perform.
4.4 Discussion
RAG for Question-Answering Over Domain-Specific Documents
The findings of this study underscore the significant role of RAG models as crucial strategies for addressing domain-specific questions. These models have shown remarkable superiority in performance compared to zero-shot knowledge and using the full PDF as context. While evaluating LLMs in well-established fields such as mathematics or biology can be straightforward, using numerous human-written sets of ground-truth questions and answers Team and Collaborators (2024), evaluating domain-specific LLMs necessitates unsupervised or semi-supervised approaches to generate evaluation benchmarks. We recognize that while our approach satisfies the need for an automated method in this domain, it still faces challenges, particularly that the selected question types might not be representative of other research areas. Therefore, researchers in other domains should carefully consider the types of questions they want to generate for their studies.
Patterns of Output
From this study, we draw two overall conclusions regarding the output patterns. First, surprisingly, we did not observe specific patterns of correctness in relation to document metadata such as token counts. This finding contradicts our initial assumption that documents with the lowest token counts would achieve the lowest accuracy and vice versa. We believe one reason for this may be that we selected only 90 passages as gold passages from the document, which might not be representative. Second, we noted that each LLM model tends to have distinctive response types. For instance, Gemini’s responses tend to be straightforward when no context is provided, often stating "I don’t have the context." In contrast, Claude and GPT-4 are more likely to attempt clarifications of input questions, such as predicting and providing the full content of abbreviations. We encourage researchers in other projects involving RAG to analyze patterns of output to enhance their work.
Long Context Reasoning
One of the primary objectives of this study is to assess how beneficial the long context models that can process 128K to 1.5M tokens context in answering questions from long EIS documents. We noticed that these models struggle to use long input contexts to answer more difficult questions that require multiple steps of reasoning (e.g, problem-solving). Given that we see the model performance varies over the positions and types of questions, we assume that effective question-type and -complexity aware reranking of retrieved documents may help to improve the performance Jeong et al. (2024). For example, we can use of another LLM to adjust the order of retrieved chunks based on the problem-solving question type.

### v2

# Examining Long-Context Large Language Models for Environmental Review Document Comprehension

@@ Abstract
As LLMs become increasingly ubiquitous, researchers have tried various techniques to augment the knowledge provided to these models. Long context and retrieval-augmented generation (RAG) are two such methods that have recently gained popularity. In this work, we examine the benefits of both of these techniques by utilizing question answering (QA) task in a niche domain. While the effectiveness of LLM-based QA systems has already been established at an acceptable level in popular domains such as trivia and literature, it has not often been established in niche domains that traditionally require specialized expertise. We construct the NEPAQuAD1.0 benchmark to evaluate the performance of five long-context LLMs – Claude Sonnet, Gemini, GPT-4, Llama 3.1, and Mistral – when answering questions originating from Environmental Impact Statements prepared by U.S. federal government agencies in accordance with the National Environmental Environmental Act (NEPA). We specifically measure the ability of LLMs to understand the nuances of legal, technical, and compliance-related information present in NEPA documents in different contextual scenarios. We test the LLMs’ internal prior NEPA knowledge by providing questions without any context, as well as assess how LLMs synthesize the contextual information present in long NEPA documents to facilitate the question/answering task. We compare the performance of the models in handling different types of questions (e.g., problem-solving, divergent, etc.). Our results suggest that RAG powered models significantly outperform those provided with only the PDF context in terms of answer accuracy, regardless of the choice of the LLM. Our further analysis reveals that many models perform better answering closed type questions (Yes/No) than divergent and problem-solving questions.
@@ 1 Introduction
As Large Language Model (LLM) s become increasingly commonplace, researchers have discovered that these models are useful for many tasks beyond text generation. Specifically, LLM s have shown potential utility in niche domains (like science) that would traditionally require specialized expertise, both in a pure text setting Horawalavithana et al. (2022); Munikoti et al. (2024) and by incorporating data of various modalities Dollar et al. (2022); Horawalavithana et al. (2023). Recent work has been done to evaluate these models Acharya et al. (2023); Munikoti et al. (2023); Cai et al. (2024) and to assess their uncertainty Wagle et al. (2024). Despite extensive research, constructing LLM s for answering domain-specific questions has proven challenging Kasneci et al. (2023).
@@
In this work, we focus on assessing the long-context LLM s in the environmental reviews conducted under the National Environmental Protection Act (NEPA)  [^1] https://www.epa.gov/nepa. NEPA is a U.S. environmental law designed to protect the environment. An environmental impact statement (EIS) is required by Section 102(2)(C) of NEPA for any proposed major federal action significantly affecting the quality of the environment. An EIS is a detailed document prepared that describes a proposed action, alternatives to the proposed action, and potential effects of the proposed action and alternatives on the environment. An EIS contains information about environmental permitting and policy decisions and considers a range of reasonable alternatives, analyzes the potential impacts resulting from the proposed action and alternatives, and demonstrates compliance with other applicable environmental laws and executive orders.
@@
In this work, we leverage both long-context LLM s and Retrieval Augmented Generation (RAG) techniques to evaluate the extent to which LLM s can comprehend NEPA knowledge for question-answering (Figure 1). We select frontier LLM s for our experiments: Claude Sonnet Anthropic Team and Collaborators (2024), Gemini Gemini Team and Collaborators (2024), GPT-4 OpenAI (2024), Llama 3.1Dubey et al. (2024) and MistralJiang et al. (2023a). We also conduct rigorous experiments evaluating LLM s with various types of contexts for NEPA documents. To evaluate our approach, we establish a benchmark using ground truth triples of questions, answers, and corresponding contexts, generated through a semi-supervised method employing GPT-4. Overall, we make the following contributions:
@@ 2.1 Gold Passage Selection
Excerpt Selection For each of nine selected documents, we attempt to select excerpts that have important content of each document. Again, the default approach of randomly extracting excerpts poses the risk of resulting in low-quality excerpts, such as parts of appendix or images’ captions. To avoid this risk, NEPA experts manually select excerpts from the documents. They divided each document into three sections: beginning, middle, and end, and then selected two, six, and two excerpts from each of these sections respectively, for a total of 10 excerpts from each document. We use these excerpts as the ground-truth context, called gold passages, for question benchmark generation.
@@ 2.2 Question Type Selection
Once we identified the gold passages, NEPA experts select the type of questions to include in the benchmark. We started with a list of 15 types of questions [^3] https://tinyurl.com/3akej8ct, and eventually narrowed it down to 10 types of questions after extensive discussions. These types are shown in Table 1. In addition to selecting the question types, the NEPA experts also created sample questions for each type. For a more detailed description of the question types, as well as example questions, please see Appendices A and B.
@@ 2.4 Ground-truth Benchmark Generation
Automated Generation We selected GPT-4 as the generative model for this task, as GPT-4 has been used for generating questions and answers for documents in other domains such as agriculture Balaguer et al. (2024). Specifically, we use GPT-4 Azure OpenAI service with default setting to execute the question generation prompts. For each document in our nine selected documents, we have 10 gold passages, results in 90 gold passages in total. For each gold passage, we generated 10 sets of closed questions and two each of the other nine types of questions. Out of these total 2520 generated questions, a significant number of these were not outputted in the correct format by the LLM s. We then filtered the generations that were not in the desired formats, bringing the number down to 1920. Finally, we manually filtered the remaining questions for quality, thus ending with a benchmark of 1599 ground-truth triples of question-answer-proof over the nine selected EIS documents. We make the benchmark publicly available. (PolicyAI, 2024).
Quality Check In order to judge the generated benchmark’s correctness, we randomly select 100 sample questions for validation. The validation was done by the NEPA experts (two co-authors). Each of them independently went through the sample questions, checking both the correctness of question type (i.e. whether the generated question was the same type of question as requested) and the correctness of the answer. For each question, if either of the evaluators marked the question or answer as correct, we labelled them as correct. Overall, our generated benchmark achieved $77\%$ of answer correctness and $82\%$ of question type correctness. While these human checks certainly helped assure the quality, we nevertheless still have some limitations in the dataset creation process, as discussed in Section 7.
@@ 3 Experimental Setup
Our experimental framework is designed to comprehensively evaluate the performance of various LLM s in responding to questions related to NEPA documents. We selected a diverse set of models, including both closed and open-source options, to provide a broad perspective on current LLM capabilities in this domain. The models under evaluation are: GPT-4 (OpenAI) OpenAI (2024), Claude-3 Sonnet (Anthropic) Anthropic Team and Collaborators (2024), Gemini (Google) Gemini Team and Collaborators (2024), 8B LLAMA3.1 (Meta)Dubey et al. (2024) and 7B Mistral (Mistral AI) Jiang et al. (2023a). To assess these models’ performance across different scenarios, we developed a benchmark consisting of questions derived from NEPA documents. Each model will be evaluated under four distinct context conditions, allowing us to analyze how different levels of information availability affect their responses. The comprehensive evaluation framework will allow us to draw meaningful conclusions about the strengths and limitations of different LLM s in handling NEPA-related queries under varying levels of contextual information.
@@ 3.1 Context Variation
We designed four distinct context scenarios that simulate various real-world information retrieval situations. These scenarios are designed to test the models’ capabilities under different levels of available context:
No Context: In this baseline scenario, questions are presented to the models without any additional context. While this strategy can work well with popular domains such as sport or literature, we assume that the NEPA domain may be challenging for the LLM s to provide accurate answers. In other words, this setting can be considered a test of the LLM s’ ability to answer out-of-general-domain questions. As such, this setting is usually expected to return low performance.
PDF Context: In this scenario, in addition to the question, we also provide the model with text contents from the entire PDF document, from which the benchmark question was generated. Since we do not inform the model which part of the document to look at, the accuracy of generated responses will heavily depend on the models’ ability to pick the correct context from the very large scale textual information provided. We expect this setting to yield performance better than the ‘no context’ scenario. This setup evaluates the models’ ability to process and extract relevant information from lengthy, complex documents.
Augmented Context:  Employing a RAG approach, this setup provides the top-$k$ most relevant text chunks retrieved from a vector database containing text chunks of NEPA documents. In our RAG model implementation, when a question is inputted for LLM generation, the corresponding context is extracted as a relevant passage from a given NEPA document. We use the standard RAG setup where we encode both question and retrieved passages with the BGE embedding model Xiao et al. (2023). Cosine similarity scores are used to assess the similarity between the question and the contexts. In our case, the number of top-ranked relevant text chunks extracted is set at $k=3$. This scenario assesses the models’ performance when given highly relevant, focused information that is specifically retrieved to answer the question at hand.
Gold Context:  In this configuration, we include the actual context from which the question was generated in the prompt, alongside the question content. This scenario represents an ideal case where the most relevant information is perfectly identified and provided. hile the situation where users manually identify the correct text passage is rare in practice, we simulate this scenario to measure how well LLM s can perform, if we were able to extract relevant passages with very high accuracy. This setup serves as an upper bound for performance, testing the models’ ability to comprehend and utilize perfectly relevant information.
@@ 3.2 Evaluation Metrics
To assess the performance of the models across different configurations, we employ a sophisticated evaluation approach that goes beyond simple syntactic similarity measures. While overlap based metrics such as BLEU scores Papineni et al. (2002) have been used in prior works, they are not ideal for evaluations where semantic accuracy is crucial. Therefore, we have chosen to use the RAGAS evaluation framework Es et al. (2023), specifically focusing on the answer correctness score. The answer correctness score combines two key aspects – (i) factual correctness: this measures the accuracy of the generated answer at the phrase/clause level compared to the ground truth. We utilize GPT-4 to quantify the factual overlap between the generated answer and the expected answer Es et al. (2023). (ii) semantic correctness: this is calculated by comparing the vector embeddings of the expected answers and the predicted answers using the BGE embedding model Xiao et al. (2023). In our evaluation, we assign different weights to these aspects – $75\%$ for factual correctness (weight of 0.75), and $25\%$ for semantic correctness (weight of 0.25). This weighting allows us to prioritize factual accuracy while still considering the overall semantic similarity of the responses. By using this comprehensive evaluation metric, we aim to provide a more nuanced and accurate assessment of the LLM s’ performance in handling NEPA-related queries across the different context scenarios. This approach enables us to capture both the precise factual content and the broader semantic accuracy of the generated responses, offering valuable insights into the models’ capabilities in processing and utilizing relevant document information.
We employ a separate evaluation method for closed type questions which have only ‘yes’ or ‘no’ answers using the following steps. First, we process the responses generated by the LLM s to extract only the ‘yes’ or ‘no’ answer from the generated response. Any additional explanatory text or elaboration is removed from the response. Thereafter, the cleaned binary response is compared directly with the ground truth answer. If the cleaned response exactly matches the ground truth, it is assigned a full score of $100\%$. If there is any discrepancy, including cases where the model fails to provide a clear "Yes" or "No" answer, the response is scored as $0\%$. This straightforward evaluation method for closed questions allows us to assess the models’ ability to provide accurate binary responses without being influenced by any uncertainties introduced by the RAGAS evaluation framework. We refer this metric as answer correctness score for closed type questions.
@@ 4 Performance Analysis
In this section, we describe the overall performance of the LLM s in the question/answering task evaluated with NEPAQuAD1.0 (as presented in Section 2). First, we compare the performance of five long-context LLM s: GPT-4, Claude-3 Sonnet, Gemini, Mistral and Llama3.1 across various QA contexts (Section 4.1). Second, we compare the model performance across various question types (Section 4.2). Finally, we analyze the performance.
@@ 4.1 Evaluating Different QA Contexts
Table 2 reports the overall performance of the models across various QA contexts used in the evaluation. We observed that for the task with no context, the GPT-4 model produces the most accurate results by far. However, when PDF documents are provided as context, Claude surpassing GPT-4 and Gemini in the answer correctness. Despite that Gemini is able to handle very long contexts (1.5M tokens), it is surprising to see its performance drop when provided with PDF documents as additional contexts. This may be due to the model struggling to reason over the large amount of relevant and irrelevant content in the EIS document.
Open-source models (LLama3.1 and Mistral) generally underperform compared to closed LLM s across most context types, including PDF, RAG, and gold contexts. However, in the no-context scenario, all models exhibit similar performance, suggesting that the availability and quality of context significantly influence the performance gap between open-source and closed models.
Overall, RAG models perform better in comparison to the models provided with PDF documents as additional contexts. In RAG setup, the Gemini model outperforms other models in term of correctness, although the correctness scores of Claude and Gemini models are much closer. There is notable increase in Llama3.1’s performance in the RAG setup when compared with the PDF contexts. This suggests that the performance of long-context models can be improved when provided with the most relevant context, as in the RAG setup.
As expected, all models perform best on average when provided with the gold passage in comparison to other context variations, as in this scenario, models synthesize information that directly contains the answer to the question posed to the model. Notably, models perform comparably when they are provided with the RAG and gold passage contexts. The only exception to this trend is the Llama3.1 model, which outperforms in the RAG setup as compared to the gold passage.
Table 2: Answer correctness of LLM responses on the NEPAQuAD benchmark across different context types. The best-performing setting for each model is shown in bold.
Context None PDF RAG Gold
GPT-4 67.00% 63.70% 74.36% 76.60%
Claude 64.53% 66.46% 75.16% 76.84%
Gemini 62.84% 65.90% 75.46% 81.15%
Mistral 64.95% 61.81% 72.88% 75.34%
Llama3.1 66.35% 59.52% 74.01% 72.73%
@@ 4.2 Evaluating Different Question Types
In this section, we summarize few interesting observations when analyzing the performance of the models across different question types. The ‘process’, ‘comparison’ and ‘evaluation’ questions generally benefit from richer contexts, with closed-source models demonstrating a more significant improvement. ‘Convergent’ and ‘recall’ type questions prove challenging for all models, especially in no-context scenarios, highlighting the difficulty of synthesizing information or retrieving specific details without supporting context. ‘Problem-solving’ questions show varied results across models, with some struggling even with rich context, indicating that complex analytical tasks remain a challenge. ‘Funnel’ questions benefit significantly from context, especially for models like Gemini 1.5, suggesting that structured, narrowing inquiries are well-suited to context-enhanced responses. The ‘closed’ questions yield the highest accuracy for all models, regardless of context. In this regard, it is important to note that the revised evaluation metric for closed questions, which directly compares the model’s ‘yes’/‘no’ response to the ground truth in contrast to the RAGAS metric used for other question types.